OpenStudy (anonymous):

1/3 log27 + 2log5 - log1

OpenStudy (anonymous):

\[=\log27^{1/3}+\log5^{2}-\log1\] \[=27^{1/3}+5^{2}-1\] \[=3+25-1\] \[=27\] I THINK.